OpenStudy (anonymous):

solve for x: 3/3x + 1/x+4 = 10/7x

Directrix (directrix):

Is this 3/(3x + 1) + 1/(x + 4) = (10/7) x ?

sam (.sam.):

(3)/(3x)+(1)/(x+4)=(10)/(7x) 1/(x)+(1)/(x+4)=(10)/(7x) Least common denominator: 7x(x+4) 14x+28=10x+40 4x+28=40 4x=12 x=3

OpenStudy (anonymous):

no its 3 over 3x + 1 over x+4 = 10 over 7x

OpenStudy (anonymous):

or 9.57 simplified

sam (.sam.):

Take my answer if 3 over 3x + 1 over x+4 = 10 over 7x